# Experience-Driven Exploration for Efficient API-Free AI Agents

### v2

@@ Abstract
Most existing software lacks accessible Application Programming Interfaces (APIs), requiring agents to operate solely through pixel-based Graphical User Interfaces (GUIs). In this API-free setting, large language model (LLM)-based agents face severe efficiency bottlenecks: limited to local visual experiences, they make myopic decisions and rely on inefficient trial-and-error, hindering both skill acquisition and long-term planning. To address these challenges, we propose KG-Agent, an experience-driven learning framework that structures an agent’s raw pixel-level interactions into a persistent State-Action Knowledge Graph (SA-KG). KG-Agent overcomes inefficient exploration by linking functionally similar but visually distinct GUI states, forming a rich neighborhood of experience that enables the agent to generalize from a diverse set of historical strategies. To support long-horizon reasoning, we design a hybrid intrinsic reward mechanism based on the graph topology, combining a state value reward for exploiting known high-value pathways with a novelty reward that encourages targeted exploration. This approach decouples strategic planning from pure discovery, allowing the agent to effectively value setup actions with delayed gratification. We evaluate KG-Agent in two complex, open-ended GUI-based decision-making environments (Civilization V and Slay the Spire), demonstrating significant improvements in exploration efficiency and strategic depth over the state-of-the-art methods.
1 College of Computer Science, Sichuan University, Chengdu, 610065, China
2 Engineering Research Center of Machine Learning and Industry Intelligence, Ministry of Education, Chengdu, 610065, China
3 Institute of High Performance Computing, Agency for Science, Technology and Research (A*STAR), Singapore 138632, Singapore
4 Centre for Frontier AI Research, Agency for Science, Technology and Research (A*STAR), Singapore 138632, Singapore
1 Introduction
@@
The emergence of Large Language Models (LLMs) has accelerated a new generation of agentic AI systems Qin et al. (2025) capable of tackling complex tasks across a diverse range of domains, including web browsing Zheng et al. (2024), operating mobile and desktop applications Zhang et al. (2025b), crafting and exploration in virtual worlds Wang et al. (2023b), and even robotics scenarios Brohan et al. (2023). Initially, the dominant paradigm relied on agents based on predefined Application Programming Interfaces (APIs), where human designers decompose high-level goals into structured workflows and tools Yao et al. (2023); Shinn et al. (2023). While this approach ensures reliability on well-defined benchmarks Xie et al. (2024), its dependence on task-specific APIs fundamentally limits the agent’s adaptability and scalability in open-ended environments. Bridging this gap, recent progress in Vision Language Models (VLMs) Zhou et al. (2022) has enabled Graphical User Interfaces (GUIs)-based agents Zhang et al. (2025a). These agents, such as UFO Zhang et al. (2024b) and CogAgent Hong et al. (2024), interact not only through APIs but also by observing and manipulating GUIs in a human-like manner. By integrating visual understanding with reasoning, they provide more general automated control and a richer, more intuitive mode of interaction. The pursuit of ultimate generality and human-like autonomy takes this progression a final step further, leading to the API-free GUI-based agents, as exemplified by CRADLE Tan et al. (2025) and Bottom-Up Agent Du et al. (2025). This methodology envisions agents that operate exclusively through a universal, human-style interface—using only screen pixels as input and keyboard/mouse actions as output. Such agents have demonstrated the remarkable ability to autonomously acquire skills from scratch in complex games and diverse software applications without any privileged access.
Despite these significant advances, as shown in Figure 1 (a), the API-free GUI-based agents face two fundamental challenges Du et al. (2025) that obstruct its path toward Artificial General Intelligence (AGI) Morris et al. (2024): (i) Inefficient Exploration: Without task-specific priors or APIs, skills must be discovered and validated purely through exhaustive trial-and-reasoning. This leads to a sample inefficiency so severe that agents typically require 2-2.5 times more environment interactions to match the progression of prior-assisted baselines. (ii) Limited Long-Horizon Strategic Reasoning: The prevailing reliance on myopic reward signals, e.g., visual changes, fails to incentivize the multi-step plans essential for sophisticated gameplay. This inability to value setup actions with delayed gratification severely limits strategic depth. Overcoming these challenges requires methods that accumulate reusable abstractions from experience and reward formulations that capture long-term value, thus enabling robust planning, transfer, and generalization across diverse tasks.
@@
Figure 1: From Local Memory to Global Strategy. (a) A conventional API-free GUI-based agent relies on a simple skill database, where experiences are isolated. Its decision-making is a reactive trial-and-error process guided by myopic rewards like visual change, leading to inefficient exploration. (b) Our proposed KG-Agent leverages a structured knowledge graph as its skill library. This graph connects experiences, enabling a strategic “exploit or explore” mechanism, and the learning of API-free GUI-based agent is guided by an advanced implicit reward that captures long-term value, accelerating knowledge transfer and fostering strategic planning.
@@
To overcome these limitations, we introduce KG-Agent, a novel experience-driven learning framework designed to transform an agent’s pixel-based GUI interactions into structured, actionable knowledge. Central to our approach is a persistent, cross-episode State-Action Knowledge Graph (SA-KG) that serves as the agent’s long-term memory. This graph organizes raw visual observations into state nodes and models acquired skills as edges between them, thereby converting unstructured pixel-level experience into a coherent network for strategic decision-making in API-free environments. To directly combat inefficient exploration, KG-Agent connects functionally analogous yet visually distinct GUI states through similarity edges, forming a rich neighborhood of experience. This enables the agent’s reasoning module to query entire functional neighborhoods, generalizing from diverse historically successful strategies rather than relying on isolated, myopic data points. To address limited long-horizon reasoning, we leverage the graph topology to design a novel hybrid intrinsic reward mechanism. This system combines a state value reward (quantifying strategic potential based on outgoing connections) with a novelty reward for environmental discovery. This potential-based formulation is crucial for incentivizing setup actions that yield delayed gratification, enabling robust long-term planning in open-ended environments. Our contribution is three-fold:
@@
  • We propose KG-Agent, a framework that structures raw GUI interaction experience into a persistent SA-KG, introducing a neighborhood of experience to combat inefficient exploration by enabling generalization across functionally similar but visually distinct states.
@@
  • We design a hybrid reward mechanism derived from the SA-KG’s topology that addresses limited long-horizon reasoning. By combining a potential-based state value reward for strategic exploitation with a novelty reward for targeted exploration, our approach effectively incentivizes multi-step plans with delayed gratification.
@@
  • We conduct extensive experiments in two complex, open-ended environments (Civilization V and Slay the Spire), demonstrating that KG-Agent significantly enhances exploration efficiency and strategic depth compared to state-of-the-art baselines.
2 Related Work
@@
LLM-based Agents. The application of LLM-based agents to complex, multi-step tasks has yielded remarkable achievements across domains Xi et al. (2025), including web browsing Gu et al. (2024), software operation Jin et al. (2024), and robotics Firoozi et al. (2025). Video games pose a particularly demanding testbed, requiring precise low-level control alongside high-level planning, abstraction, and adaptation Li et al. (2025b). Recent advances have shown strong performance in environments like Minecraft Li et al. (2025a) and StarCraft II Ma et al. (2024), yet early work often relied on privileged APIs for simplified observations and semantic action spaces Wang et al. (2023a), limiting applicability to closed-source commercial games that expose only raw pixels. With the rise of multimodal capabilities, GUI-based agents have emerged Zhang et al. (2025a), such as OpenAI Operator, UFO Zhang et al. (2024b), and CogAgent Hong et al. (2024), which integrate visual understanding with text-based reasoning to enable more general and intuitive control of diverse applications.
@@
API-free Agents. In pursuit of the ultimate, human-like General AI, a new paradigm of agents that operate purely from raw pixel inputs without any API support has emerged. Exemplified by works like CRADLE Tan et al. (2025) and the Bottom-Up Agent Du et al. (2025), these systems learn complex skills from scratch via a universal human-style interface (screen, keyboard, and mouse), using an autonomous trial-and-error loop to discover and refine behaviors. This API-free GUI-based approach has proven capable of achieving tangible progress in complex open-ended environments. However, this promising paradigm introduces a critical challenge: experience siloing. Agents often retrieve memory in a localized and myopic way—relying on the most visually similar past state—failing to generalize across functionally similar but visually different contexts. As a result, they fall back on inefficient trial-and-error relearning, hindering knowledge accumulation and long-term strategy formation.
3 Methodology
@@
As shown in Figure 2, the proposed KG-Agent is composed of three core components: a universal Environment IO Interface for open-ended environment interaction, a Memory System to store and structure experience, and a VLM-based Reasoning Module that directs the agent’s behavior. Following Bottom-Up Agent Du et al. (2025), we model the environment as a Partially Observable Markov Decision Process (POMDP) Spaan (2012) defined by $(\mathcal{X},\mathcal{A},\mathcal{T},\mathcal{R})$, where $\mathcal{X}$ is the visual observation space, $\mathcal{A}$ is the set of atomic actions, $\mathcal{T}$ represents the unknown transition dynamics, and $\mathcal{R}$ is the implicit reward signal. The agent is built around a VLM $f_{VLM}$ that, when conditioned on a prompt promptprompt and current context $c_{i}$, produces function-specific outputs: $f_{VLM}(prompt,c_{i})\rightarrow y_{i}$. A skill $\sigma$ is defined as a sequence of atomic actions $\sigma=(a_{1},…,a_{k})$, each paired with a semantic descriptor $d_{\sigma}$, a natural language summary of its intent, generated by the VLM. The skill library $\mathbb{S}=\{\sigma_{1},…,\sigma_{n}\}$ evolves over time through discovery, refinement, and composition.
Figure 2: Overview of the proposed KG-Agent. For a given state $i$, the agent queries its memory to find a neighborhood of experience ($\mathcal{N}_{\mathcal{E}}$), where state nodes are connected by similarity edges (blue line) or skill edges (red arrow). If candidate skills $\mathbb{S}_{i}$ are retrieved, they are executed to transition the environment to state $j$. In addition, a VLM is central to this agent’s reasoning, performing skill invocation, refinement, augmentation, and evaluation, where the refinement is guided by a hybrid implicit reward composed of state value reward and novelty reward.
3.1 The Environment IO Interface
@@
To operate in API-free environments, our agent requires a universal interface for perception and action. The Environment IO Interface fulfills this role, serving as the sole conduit through which the agent perceives and interacts with its environment. It operates exclusively on raw visual input and produces low-level, human-like actions, without access to internal states or privileged APIs. This design ensures our agent’s generalizability while posing significant challenges in visual understanding and precise action execution.
@@
Environment Input: Purely Visual Observations. The agent’s perception is grounded exclusively in raw visual observations. At each time step $i$, the agent receives a screenshot $X_{i}\in\mathcal{X}$, which constitutes its entire sensory input. From this screenshot, a feature vector $x_{i}$ is extracted using the image encoder of a pre-trained CLIP model Radford et al. (2021) to represent the state. Notably, our agent forgoes the use of any Optical Character Recognition (OCR) models to extract textual content. This forces the agent to rely entirely on the VLM’s intrinsic spatial perception and visual understanding to interpret the scene, without access to structured information.
@@
Environment Output: Simulated Mouse and Keyboard Operations. The agent interacts with the environment by generating low-level keyboard and mouse commands from the action space $\mathcal{A}$, mirroring human interaction. The action space includes all possible operations such as click, drag, scroll, and type. These can be combined in various ways to form combos and shortcuts. Considering that mouse operations can theoretically target any pixel on the screen, making the action space excessively large, we employ the Segment Anything Model (SAM) Kirillov et al. (2023) to dynamically identify and segment interactable UI elements from the current observation $X_{i}$. The interactable UI elements are also updated in the memory. A skill $\sigma$ is typically composed of one or more actions parameterized by the objects of interaction. For example, the skill ’Choose Production and Build Monument’ translates to the operations: {’operate’: ’Click’, ’object_id’: 11, ’object_name’: ’Choose_Production’}; {’operate’: ’Click’, ’object_id’: 24, ’object_name’: ’Monument’}. To execute such an operation, the agent first retrieves a reference image of the target object (e.g., the ’Choose_Production’ button) from memory. It then performs template matching against the current screen $X_{i}$ to find the object’s precise location, calculating the center coordinates for the mouse click. This process grounds the VLM’s symbolic action plan into concrete, executable operating system-level keyboard and mouse operations on the GUI.
3.2 Experience-Driven Memory System
@@
The memory of KG-Agent is designed to store and manage all useful information, enabling the agent to learn from past experiences and make informed decisions. It consists of two complementary components: a Procedural Memory and a SA-KG.
@@
Procedural Memory. This component maintains the agent’s immediate, operational knowledge, including executable skills and cached plans. It is implemented as a set of structured tables: i) Objects Table: Stores information about interactable UI elements identified by SAM, including their names and reference images. ii) Skill Table: Contains the evolving skill library $\mathbb{S}$, where each skill $\sigma$ is defined by its name, semantic descriptor $d_{\sigma}$, a sequence of operations $\{a_{1},a_{2},...,a_{k}\}$, and a fitness score $\phi_{\sigma}$ that reflects its historical effectiveness. iii) Action Clusters Table: Groups skills that are applicable in similar states (identified by their feature vectors), facilitating efficient retrieval of contextually relevant skills. iv) Monte Carlo Search Trees (MCST) Table: Stores serialized search trees Browne et al. (2012) associated with specific states, allowing the agent to cache and reuse complex decision-making plans.
@@
State-Action Knowledge Graph (SA-KG). The SA-KG is the cornerstone of the agent’s long-term memory, designed specifically to overcome the bottleneck of inefficient exploration. It is a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ that transforms the agent’s memory from a passive store of episodes into a semantically connected network that facilitates strategic reasoning and generalization. The SA-KG is constructed from nodes $\mathcal{V}$ and two types of edges: similarity edges $\mathcal{E}^{Sim}$ and skill edges $\mathcal{E}^{\sigma}$: i) Nodes ($\mathcal{V}$): Each node $v_{i}\in\mathcal{V}$ represents a unique state, characterized by its CLIP feature vector $x_{i}$. A new observation with feature vector $x_{j}$ is merged into an existing node $v_{i}$ if their cosine similarity exceeds a threshold, i.e., $\cos(x_{i},x_{j})>\theta_{{merge}}$; otherwise, a new node $v_{j}$ is created. ii) Similarity Edges ($\mathcal{E}^{Sim}$): An undirected edge $e^{\mathrm{sim}}_{i,j}$ connects two functionally analogous but visually distinct states $v_{i}$ and $v_{j}$ if $\theta_{{merge}}>\cos(x_{i},x_{j})>\theta_{{simi}}$. Each edge has a weight $w^{\mathrm{sim}}_{i,j}\in[0,1]$ equal to the cosine similarity between $x_{i}$ and $x_{j}$. iii) Skill Edges ($\mathcal{E}^{\sigma}$): A directed edge $e^{\sigma}_{i,j}$ from state $v_{i}$ to $v_{j}$ represents the successful execution of a skill $\sigma$. Each skill edge has a weight $w^{\sigma}_{i,j}$ quantifying the utility of that skill $\sigma$ for the specific state transition. Crucially, some actions produce a large immediate visual change but lead to a dead end, while a crucial setup move might have minimal impact but be key to a long-term strategy. To capture this, we define the skill edge weight $w^{\sigma}_{i,j}$ as a combination of immediate visual change and historical effectiveness:
@@
where $f_{sigmoid}(\cdot)$ is the Sigmoid activation function, the $\Delta_{i,j}$ is the visual change ratio between states $v_{i}$ and $v_{j}$, the $\phi_{\sigma}$ is the historical fitness of skill $\sigma$, the $C_{0}$ is a constant that controls the sensitivity of the fitness term, and the $\alpha$ is a weighting factor that balances the importance of immediate change versus long-term strategic value. This design allows the agent to develop sophisticated, long-horizon plans by valuing both types of actions.
3.3 Reasoning and Decision-Making Module
The reasoning module is the decision-making engine of KG-Agent, responsible for a continuous cycle of skill invocation, augmentation, refinement, and evaluation. It orchestrates a hierarchical decision-making process that leverages the structured knowledge in both the SA-KG and Procedural Memory to intelligently balance exploiting known strategies with exploring new possibilities. This entire process is guided by a novel, graph-based hybrid reward mechanism.
Skill Invocation. The agent employs a hierarchical, two-stage process for skill invocation that prioritizes structured, experience-based knowledge from the SA-KG before falling back to a more general VLM-guided search, ensuring both efficiency and reliability. Upon entering a new state $v_{i}$, the agent first identifies its Neighborhood of Experience $\mathcal{N}_{\mathcal{E}}(v_{i})$, defined as the set of all nodes connected to $v_{i}$ by similarity edges:

### v3

# SAG-Agent: Enabling Long-Horizon Reasoning in Strategy Games via Dynamic Knowledge Graphs

@@ Abstract
Most commodity software lacks accessible Application Programming Interfaces (APIs), requiring autonomous agents to interact solely through pixel-based Graphical User Interfaces (GUIs). In this API-free setting, large language model (LLM)-based agents face severe efficiency bottlenecks: limited to local visual experiences, they make myopic decisions and rely on inefficient trial-and-error, hindering both skill acquisition and long-horizon planning. To overcome these limitations, we propose SAG-Agent, an experience-driven learning framework that structures an agent’s raw pixel-level interactions into a persistent State-Action Graph (SAG). SAG-Agent mitigates inefficient exploration by topologically linking functionally similar but visually distinct GUI states, constructing a rich neighborhood of experience that enables the agent to generalize from a diverse set of historical strategies. To facilitate long-horizon reasoning, we design a novel hybrid intrinsic reward mechanism based on the graph topology, combining a state-value reward for exploiting known high-value pathways with a novelty reward that encourages targeted exploration. This approach decouples strategic planning from pure discovery, allowing the agent to effectively value setup actions with delayed gratification. We evaluate SAG-Agent in two complex, open-ended GUI-based decision-making environments (Civilization V and Slay the Spire), demonstrating significant improvements in exploration efficiency and strategic depth over the state-of-the-art methods.
I Introduction
@@
The emergence of Large Language Models (LLMs) has accelerated a new generation of agentic Artificial Intelligence (AI) systems [1] capable of tackling complex tasks across a diverse range of domains, including web browsing [2], operating mobile and desktop applications [3], crafting and exploration in virtual worlds [4], and even robotics scenarios [5]. Initially, the dominant paradigm relied on agents based on predefined Application Programming Interfaces (APIs), where human designers decompose high-level goals into structured workflows and tools [6]. While this approach ensures reliability on well-defined benchmarks [7], its dependence on task-specific APIs fundamentally limits the agent’s adaptability and scalability in open-ended environments. Bridging this gap, recent progress in Vision Language Models (VLMs) [8] has enabled Graphical User Interfaces (GUIs)-based agents [9]. These agents, such as UFO [10] and CogAgent [11], interact not only through APIs but also by observing and manipulating GUIs in a human-like manner. By integrating visual understanding with reasoning, they provide more general automated control and a richer, more intuitive mode of interaction. The pursuit of ultimate generality and human-like autonomy takes this progression a final step further, leading to the API-free GUI-based agents, as exemplified by CRADLE [12] and Bottom-Up Agent [13]. This methodology envisions agents that operate exclusively through a universal, human-style interface—using only screen pixels as input and keyboard/mouse actions as output. Such agents have demonstrated the remarkable ability to autonomously acquire skills from scratch in complex games and diverse software applications without any privileged access.
@@
Figure 1: From Local Memory to Global Strategy. (a) A conventional API-free GUI-based agent relies on a simple skill database, where experiences are isolated. Its decision-making is a reactive trial-and-error process guided by myopic rewards, leading to inefficient exploration. (b) Our agent leverages a structured knowledge graph as its skill library. This graph connects experiences, enabling a strategic exploit or explore mechanism, and the learning of agent is guided by an advanced implicit reward that captures long-term value, accelerating knowledge transfer and fostering strategic planning.
Despite these significant advances, as shown in Fig. 1 (a), the API-free GUI-based agents face two fundamental challenges [13, 12] that obstruct their path toward Artificial General Intelligence (AGI) [14]: (i) Inefficient Exploration: Without task-specific priors or APIs, skills must be discovered and validated purely through exhaustive trial-and-reasoning. This leads to a sample inefficiency so severe that agents typically require 2-2.5 times more environment interactions to match the progression of prior-assisted baselines. (ii) Limited Long-Horizon Strategic Reasoning: The prevailing reliance on myopic reward signals, e.g., visual changes, fails to incentivize the multi-step plans essential for sophisticated gameplay. This inability to value setup actions with delayed gratification severely limits strategic depth. Overcoming these challenges requires methods that accumulate reusable abstractions from experience and reward formulations that capture long-term value, thus enabling robust planning, transfer, and generalization across diverse tasks.
@@
To overcome these limitations, we introduce SAG-Agent, a novel experience-driven learning framework designed to transform an agent’s pixel-based GUI interactions into structured, actionable knowledge. Central to our approach is a persistent, cross-episode State-Action Graph (SAG) that serves as the agent’s long-term memory. This graph organizes raw visual observations into state nodes and models acquired skills as edges between them, thereby converting unstructured pixel-level experience into a coherent network for strategic decision-making in API-free environments. To directly combat inefficient exploration, SAG-Agent connects functionally analogous yet visually distinct GUI states through similarity edges, forming a rich neighborhood of experience. This enables the agent’s reasoning module to query entire functional neighborhoods, generalizing from diverse historically successful strategies rather than relying on isolated, myopic data points. To address limited long-horizon reasoning, we leverage the graph topology to design a novel hybrid intrinsic reward mechanism. This system combines a state value reward (quantifying strategic potential based on outgoing connections) with a novelty reward for environmental discovery. This potential-based formulation is crucial for incentivizing setup actions that yield delayed gratification, enabling robust long-term planning in open-ended environments. Our contribution is three-fold:
@@
  • We propose SAG-Agent, a framework that structures raw GUI interaction experience into a persistent SAG, introducing a neighborhood of experience to combat inefficient exploration by enabling generalization across functionally similar but visually distinct states.
@@
  • We design a hybrid reward mechanism derived from the SAG’s topology that addresses limited long-horizon reasoning. By combining a potential-based state value reward for strategic exploitation with a novelty reward for targeted exploration, our approach effectively incentivizes multi-step plans with delayed gratification.
@@
  • We conduct extensive experiments in two complex, open-ended environments (Civilization V and Slay the Spire), demonstrating that SAG-Agent significantly enhances exploration efficiency and strategic depth compared to state-of-the-art baselines.
II Related Work
@@
LLM-based Agents. The application of LLM-based agents to complex, multi-step tasks has yielded remarkable achievements across diverse domains [15], including web browsing [16], software operation [17], and robotics [18]. Video games pose a particularly demanding testbed, requiring precise low-level control alongside high-level planning, abstraction, and adaptation [19]. Recent advances have shown strong performance in environments like Minecraft [20] and StarCraft II [21], yet early work often relied on privileged APIs for simplified observations and semantic action spaces [22], limiting applicability to closed-source commercial games that expose only raw pixels. Concurrently, the integration of LLMs [23] into multi-agent systems has established a new architectural paradigm in which agents actively reason, communicate, and interact with external tools, environments, and users through structured interfaces [24]. This paradigm elevates LLMs from passive text generators to dynamic participants capable of orchestrating complex tasks [25]. Notably, the rise of multimodal capabilities has spurred the development of GUI-based agents [9], e.g., OpenAI Operator, UFO [10], and CogAgent [11], which integrate visual understanding with text-based reasoning to enable general and intuitive control. In practice, modern LLM-based agents function as autonomous or semi-autonomous entities that combine natural language understanding with tool-augmented reasoning [26], with GUIs further extending their interactive capacity [27]. In simulated or game environments, GUI perception modules can translate visual elements (e.g., buttons, maps, menus) into semantic representations for LLM reasoning [28], enabling operation across diverse digital environments without task-specific policy training [29]. Frameworks such as AutoGen [30], MetaAgent [31], and Generative Agents [32] formalize these patterns, offering abstractions for agent roles, tool integration, message routing, and environment interfacing. By unifying natural language communication with API-driven and GUI-mediated interaction, LLM-based multi-agent systems are advancing from conceptual models to deployable platforms for open-ended, human-aligned reasoning and decision-making.
@@
API-Free LLM-based Agents. While early LLM-based agents frequently depended on privileged access to internal game states via APIs, this API-centric paradigm faces critical constraints [22]. Its reliance on engineered interfaces inherently limits generalization, particularly in closed-source commercial games and proprietary software where internal APIs are inaccessible or undocumented [12], and distances the agent from the rich visual context of genuine human-computer interaction. In pursuit of ultimate generality, a new paradigm has emerged: API-free LLM-based agents that operate purely from raw pixel input, interacting through universal human-style interfaces without relying on internal state APIs [12, 13]. Exemplified by systems like CRADLE [12] and the Bottom-Up Agent [13], these agents learn complex skills from scratch via an autonomous trial-and-error loop. This approach has demonstrated tangible progress in open-ended environments, such as multimodal desktop assistants (e.g., UFO [10], CogAgent [11]) and game-playing agents, achieving meaningful in-game progress without game-specific adaptations. However, this promising paradigm introduces a critical challenge: experience siloing. Agents often retrieve memory in a localized and myopic way, relying on the most visually similar past state, and consequently fail to generalize across functionally similar but visually different contexts. This forces them to fall back on inefficient trial-and-error relearning, which hinders systematic knowledge accumulation and long-term strategy formation, underscoring a key obstacle to achieving truly general and flexible AI systems.
III Methodology
@@
As shown in Fig. 2, the proposed SAG-Agent is composed of an Environment IO Interface for environment interaction, a Memory System to store and structure experience, and a VLM-based Reasoning Module that directs the agent’s behavior. Following Bottom-Up Agent [13], we model the environment as a Partially Observable Markov Decision Process [33] defined by $(\mathcal{X},\mathcal{A},\mathcal{T},\mathcal{R})$, where $\mathcal{X}$ is the visual observation space, $\mathcal{A}$ is the set of atomic actions, $\mathcal{T}$ represents the unknown transition dynamics, and $\mathcal{R}$ is the implicit reward signal. The agent is built around a VLM $f_{*}$ that, when conditioned on a prompt $\mathbb{P}$ and current context $c_{i}$, produces function-specific outputs: $f_{*}(\mathbb{P},c_{i})\rightarrow y_{i}$. A skill $\sigma$ is defined as a sequence of atomic actions $\sigma=(a_{1},…,a_{k})$, each paired with a semantic descriptor $d_{\sigma}$, a natural language summary of its intent, generated by the VLM. The skill library $\mathbb{S}=\{\sigma_{1},…,\sigma_{n}\}$ evolves over time through discovery, refinement, and composition.
III-A The Environment IO Interface
@@
To operate in API-free environments, our agent requires a universal interface for perception and action. The Environment IO Interface, fulfills this role, serving as the sole conduit through which the agent perceives and interacts with its environment. It operates exclusively on raw visual input and produces low-level, human-like actions, without access to internal states or privileged APIs. This design ensures our agent’s generalizability while posing significant challenges in visual understanding and precise action execution.
@@
Environment Input: Purely Visual Observations. The agent’s perception is grounded exclusively in raw visual observations. At each time step $i$, the agent receives a screenshot $X_{i}\in\mathcal{X}$, which constitutes its entire sensory input. From this screenshot, a feature vector $x_{i}$ is extracted using the image encoder of a pre-trained CLIP model [34] to represent the state. Notably, our agent forgoes the use of any Optical Character Recognition (OCR) models to extract textual content. This forces the agent to rely entirely on the VLM’s intrinsic spatial perception and visual understanding to interpret the scene, without access to structured information.
@@
Environment Output: Simulated Mouse and Keyboard Operations. The agent interacts with the environment by generating low-level keyboard and mouse commands from the action space $\mathcal{A}$, mirroring human interaction. The action space includes all possible operations such as click, drag, scroll, and type. These can be combined in various ways to form combos and shortcuts. Considering that mouse operations can theoretically target any pixel on the screen, making the action space excessively large, we employ the Segment Anything Model (SAM) [35] to dynamically identify and segment interactable UI elements from the current observation $X_{i}$. The interactable UI elements are also updated in the memory. A skill $\sigma$ is typically composed of one or more actions parameterized by the objects of interaction. For example, the skill ’Choose Production and Build Monument’ translates to the operations: {’operate’: ’Click’, ’object_id’: 11, ’object_name’: ’Choose_Production’}; {’operate’: ’Click’, ’object_id’: 24, ’object_name’: ’Monument’}. To execute such an operation, the agent first retrieves a reference image of the target object (e.g., the ’Choose_Production’ button) from memory. It then performs template matching against the current screen $X_{i}$ to find the object’s precise location, calculating the center coordinates for the mouse click. This process grounds the VLM’s symbolic action plan into concrete, executable operating system-level keyboard and mouse operations on the GUI.
Figure 2: Overview of the proposed SAG-Agent. For a given state $i$, the agent queries its memory to find a neighborhood of experience ($\mathcal{N}_{\mathcal{E}}$), where state nodes are connected by similarity edges (blue line) or skill edges (red arrow). If candidate skills $\mathbb{S}_{i}$ are retrieved, they are executed to transition the environment to state $j$. In addition, a VLM is central to this agent’s reasoning, performing skill invocation, refinement, augmentation, and evaluation.
III-B Experience-Driven Memory System
@@
The memory of SAG-Agent is designed to store and manage all useful information, enabling the agent to learn from past experiences and make informed decisions. It consists of two components: a Procedural Memory and an SAG.
@@
Procedural Memory. Inspired by Bottom-Up Agent [13], this component maintains the agent’s immediate, operational knowledge, including executable skills and cached plans. It is implemented as a set of structured tables: i) Objects Table: Stores information about interactable UI elements identified by SAM, including their names and reference images. ii) Skill Table: Contains the evolving skill library $\mathbb{S}$, where each skill $\sigma$ is defined by its name, semantic descriptor $d_{\sigma}$, a sequence of operations $\{a_{1},a_{2},...,a_{k}\}$, and a fitness score $\phi_{\sigma}$ that reflects its historical effectiveness. iii) Action Clusters Table: Groups skills that are applicable in similar states (identified by their feature vectors), facilitating efficient retrieval of contextually relevant skills. iv) Monte Carlo Tree Search (MCTS) Table: Stores serialized search trees [36] associated with specific states, allowing the agent to cache and reuse complex decision-making plans.
@@
State-Action Graph (SAG). The SAG is the cornerstone of the agent’s long-term memory, designed specifically to overcome the bottleneck of inefficient exploration. It is a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ that transforms the agent’s memory from a passive store of episodes into a semantically connected network that facilitates strategic reasoning and generalization. The SAG is constructed from nodes $\mathcal{V}$ and two types of edges: similarity edges $\mathcal{E}^{Sim}$ and skill edges $\mathcal{E}^{\sigma}$: i) Nodes ($\mathcal{V}$): Each node $v_{i}\in\mathcal{V}$ represents a unique state, characterized by its CLIP feature vector $x_{i}$. A new observation with feature vector $x_{j}$ is merged into an existing node $v_{i}$ if their cosine similarity exceeds a threshold, i.e., $\cos(x_{i},x_{j})>\theta_{{merge}}$; otherwise, a new node $v_{j}$ is created. ii) Similarity Edges ($\mathcal{E}^{Sim}$): An undirected edge $e^{\mathrm{sim}}_{i,j}$ connects two functionally analogous but visually distinct states $v_{i}$ and $v_{j}$ if $\theta_{{merge}}>\cos(x_{i},x_{j})>\theta_{{simi}}$. Each edge has a weight $w^{\mathrm{sim}}_{i,j}\in[0,1]$ equal to the cosine similarity between $x_{i}$ and $x_{j}$. iii) Skill Edges ($\mathcal{E}^{\sigma}$): A directed edge $e^{\sigma}_{i,j}$ from state $v_{i}$ to $v_{j}$ represents the successful execution of a skill $\sigma$. Each skill edge has a weight $w^{\sigma}_{i,j}$ quantifying the utility of that skill $\sigma$ for the specific state transition. Crucially, some actions produce a large immediate visual change but lead to a dead end, while a crucial setup move might have minimal impact but be key to a long-term strategy. To capture this, we define the skill edge weight $w^{\sigma}_{i,j}$ as a combination of immediate visual change and historical effectiveness:
@@
where $f_{sigmoid}(\cdot)$ is the Sigmoid activation function, the $\Delta_{i,j}$ is the visual change ratio between states $v_{i}$ and $v_{j}$, the $\phi_{\sigma}$ is the historical fitness of skill $\sigma$, the $C_{0}$ is a constant that controls the sensitivity of the fitness term, and the $\alpha$ is a weighting factor that balances the importance of immediate change versus long-term strategic value. We quantify visual change as the proportion of changed pixels between consecutive grayscale screenshots. Pixels with an absolute difference exceeding 30 (0–255 scale) are counted, and the ratio $\Delta_{i,j}$ of such pixels to total pixels is computed. This design allows the agent to develop sophisticated, long-horizon plans by valuing both types of actions.
III-C Reasoning and Decision-Making Module
The reasoning module is the decision-making engine of SAG-Agent, responsible for a continuous cycle of skill invocation, augmentation, refinement, and evaluation. It orchestrates a hierarchical decision-making process that leverages the structured knowledge in both the SAG and Procedural Memory to intelligently balance exploiting known strategies with exploring new possibilities. This entire process is guided by a novel, graph-based hybrid reward mechanism.
Skill Invocation. The SAG-Agent employs a hierarchical, two-stage process for skill invocation that prioritizes structured, experience-based knowledge from the SAG before falling back to a more general VLM-guided search, ensuring both efficiency and reliability. Upon entering a new state $v_{i}$, the agent first identifies its Neighborhood of Experience $\mathcal{N}_{\mathcal{E}}(v_{i})$, defined as the set of all nodes connected to $v_{i}$ by similarity edges: